OpenStudy (anonymous):

check my sisters answers please what is the slope of the line that passes through the pair of points (-7/2,-3) and (-5, 5/2)? 6/22 -6/22 22/6 -22/6

OpenStudy (anonymous):

she says she thinks it is 22/6 is she right?

OpenStudy (anonymous):

y^2-y^1 ------- x^2-x^1 All you have to do is plug the points into the equation

OpenStudy (anonymous):

@emilyhaddad

OpenStudy (anonymous):

it's -22/6

OpenStudy (anonymous):

ok ill explain the signs to her she forgot that they switch sometimes during the whole transition of it all

OpenStudy (anonymous):

also she wants to know if this one is right write an equation in point slope form for the line through the given point with the given slope (5, -6); m = -0.38 y + 6 = -0.38(x + 5) y - 6 = -0.38x - 5 y - 6 = -0.38(x + 5) y + 6 = -0.38(x - 5) she thinks it is the last one is she right?

OpenStudy (anonymous):

Yes sir!

OpenStudy (anonymous):

hah thx she really appreciates it

OpenStudy (anonymous):

no problem!

OpenStudy (anonymous):

hey she has another question willing to help?

OpenStudy (anonymous):

for the data in the table does y vary directly with x if it does write an equation for the direct variation x y 32 24 16 8 8 6 yes y = 0.5x yes y = 2x yes y = 1.5x no y does not vary directly with x she says she thinks it is yes y = 2x

OpenStudy (anonymous):

@Emmalineh123